OpenStudy (anonymous):

f(x)=4x-3 g(x)=x^3+2x Find (f-x)(4) Subject: Algebra 2

OpenStudy (anonymous):

@dan815

OpenStudy (shadowlegendx):

@dan815 This user was calling for your help and you IGNORED HIM?!?!?!?!?!?!? I abhor you dan, you asinine pig, so malevolent to these students that just want to learn from your intellect but you selfishly keep it all to your self.

OpenStudy (dan815):

f(x)=4x-3 g(x)=x^3+2x Find (f-g*)(4)

OpenStudy (dan815):

(f-g)(X) = 4x-3 - (x^3+2x) = -x^3+2x-3 (f-g)(4)=-x^3+2x-3 =-(4)^3+2(4)-3=-64+8-3=-59

OpenStudy (shadowlegendx):

*claps*

OpenStudy (shadowlegendx):

@dan815 I forgive you

OpenStudy (shadowlegendx):

You sins have been rectified and forgiven before the eyes of God